# Fading in the Flow: Suppression of cold gas growth in expanding galactic outflows

### v2

@@ 3.3 Properties of clouds in a typical CC85 wind
Figure 6: Temporal evolution of the profile of cloud ($T<2T_{\rm cl}=8\times 10^{4}\ \rm K$) pressure (left panel), density (middle panel), and velocity (right panel). The median cloud profiles are shown in thin black bordered lines, while all the cloud material is shown by dots at times 0 (red), 15 (pink), 45 (yellow) $t/t_{\rm cc,ini}$. The cloud profiles are normalized by the wind properties at the initial location of the cloud. Distance down the wind is expressed in multiples of the initial position of the cloud (w.r.t. the wind center), and also in kpc (top labels). The simulation chosen for this plot corresponds to our fiducial parameters $(\mathcal{M},\chi,t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini})=(1.496,100,0.2)$; cf. Table 1. The black solid lines in all the panels show the background wind profiles, and the vertical lines mark the instantaneous position of the center of mass of the cold cloud (with the same color as their profile). The left panel demonstrates that the cold cloud always maintains an approximate local pressure equilibrium with the wind. This implies that the cloud density is proportional to the wind pressure ($n_{\rm cl}\propto d_{\rm cl}^{-2\gamma}$; cf. Eq. 3.1) at any location (because the cloud temperature is essentially fixed; cf. Fig. 8) indicated by the thin gray dotted line in the middle panel. The cloud median density profile approximately follows this gray line. The cloud velocity profile on the right panel shows that it gradually attains the wind velocity, with its tail becoming co-moving faster than the head. In addition, the black dashed line in the middle panel shows the profile of $10$ times the local wind density for reference as this denotes the density cutoff used in our work to identify clouds that stand out as contrasting features with respect to their background (see Section 3.4 & Fig. 7 for details). Animated versions of this plot are available in the following links for the profiles of pressure, density, and velocity.
@@
In each snapshot from our simulation, starting from the minimum until the maximum radial extent of cold gas along the wind, we choose concentric radial shells in the wind of width $0.25\ R_{\rm cl}$. We calculate the median pressure (mass-weighted), density, and velocity of the cold gas ($T<8\times 10^{4}\ \rm K$) in each of these shells to obtain a cloud profile as a function of the downstream distance (for fiducial parameters; cf. Table 1). These pressure, density, and velocity profiles of the cloud are presented in Fig. 6, where we also show all the cloud cells as an underlying scatter plot to indicate the spread in the data. We use black dashed lines to show the corresponding CC85 wind profile for reference. Different colors are chosen to denote cloud profiles at different times in their evolution, as labeled in the accompanying legends. From these cloud profiles, we find that the cloud pressure, in the left panel of Fig. 6, closely follows the local wind pressure (except for the small portion at the head of the cloud). Therefore, the cloud evolves roughly isobarically in the wind. Note that the median pressure is slightly lower than the local wind pressure which might be reponsible for the inflow of cold gas into the cloud as our fiducial parameter corresponds to a growing cloud with strong radiative cooling (see Section 3.5 & Fig. 8 for details). In addition, since the cold gas itself is approximately isothermal (see Fig. 4 & Section 3.5), we can approximate the cloud density as
@@
which implies that $n_{\rm cl}\propto p_{\rm wind}$. Therefore, in the asymptotic limit, $n_{\rm cl}\propto(d_{\rm cl}/d_{\rm cl,ini})^{-2\gamma}$, where $d_{\rm cl}$ and $d_{\rm ini}$ are the distances of the cold gas and the initial position of the cloud respectively from the wind center. Indeed the cloud density profile follows our expectation as indicated by the dotted gray curves in the middle panel of Fig. 6. The most important implication is that the cloud loses its density contrast with distance. This is generic for growing clouds of all sizes, causing them to gradually dissolve in the background wind (cf. Section 3.4). The cloud velocity profile in the rightmost panel in Fig. 6 shows that the cloud gradually catches up with the velocity of the wind. As the wind velocity in the CC85 solution is not strictly a constant (but asymptotes to a constant), there is a substantial velocity difference between the head and the tail of the cloud, causing it to stretch in the direction of the wind.
In addition to this, we find that the inflow velocity of gas due to radiative cooling feeding the cold clouds as it moves in the CC85 wind shows only a crude agreement with the theory of turbulent boundary layers introduced by Fielding et al. 2020; Gronke & Oh 2020; Tan et al. 2021 (see Appendix B). Further work is needed to test the limits of applicability of this theory and to come up with better models.
@@ 3.4 Clouds as dense cold clumps embedded in a diffuse background
Figure 7: Temporal evolution of mass in clouds, defined as cold gas ($T<2T_{\rm cl}=8\times 10^{4}\ \rm K$) much denser than their local background (more than $10$ times denser than the wind at their location). The gas cells satisfying this criterion stand out as contrasting radial filaments in the wind (cf. Fig. 4). The time is normalized by the initial cloud-crushing time $t_{\rm cc,ini}$ (which is larger for larger and vice versa). Each line in the plot refers to a simulation with a fixed value of $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}$ (inversely proportional to the initial cloud size) – annotated by the numbers adjacent to the curves. The lines are colored according to the distance the cloud has travelled with respect to its initial position in the wind. Since the Mach number at the initial position $\mathcal{M}=1.496$ of the cloud and its density contrast $\chi=100$ are kept the same across all the simulations presented here, the initial cloud position $d_{\rm cl,ini}$ is the same. The cloud density falls at a steeper rate with distance compared to the the CC85 wind density (cf. Section 3.3). Consequently, the clouds fail to remain cold as dense contrasting features in the wind as they travel downstream. Therefore, after the cloud has travelled a critical distance down the wind (approximately the same in all simulations; cf. Eq. 3.1), the (contrasting) cloud mass transitions from an initially growing phase to a destruction phase. This transition occurs at around $d_{\rm cl}/d_{\rm cl,ini}\sim 4$, where the line color is blue-cyan and the gray dotted and black dashed lines intersect in the middle panel of Fig. 6.
@@
In the context of cloud-crushing in a CC85 wind, the cloud encounters a cooler adiabatically expanding wind as it moves downstream. Consequently, the cold cloud formed by mixing and subsequent radiative cooling becomes diffuse as it moves further downstream. Therefore, simply identifying all the cold gas as cloud material gives an incomplete picture of the ongoing process. Considering this, we expand the definition of cloud as features formed (due to mixing and radiative cooling) as cold and dense (not just cold) relative to the featureless expanding wind. (cf. Section 2.3). In the remainder of this section we illustrate that in the presence of a CC85 wind, the clouds in the cooling regime grow in their cold mass but at a reduced rate (see Section 3.2 & Fig. 5), and in addition continue to expand compared to clouds in a uniform plane-parallel wind. It is worth noting that the ratio of the contrast (with the background) that a cloud has between any two distances down the wind is the same for clouds of all sizes (cf. Section 3.3) if the clouds start from the same initial position with the same contrast. This ensures that the qualitative picture that clouds lose contrast while moving downstream remains unaltered by the choice of threshold density contrast ($10$ here; see Section 2.3).
@@
We illustrate the evolution of the cloud mass identified using our ‘contrasting cloud identification’ criterion (discussed in Section 2.3) in Fig. 7. The different curves in the figure correspond to different initial cloud sizes in our simulation, labeled by $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}$. The clouds in different runs from our simulation presented here are initially placed at the same position with respect to the wind (for same Mach number $\mathcal{M}_{\rm ini}$) with identical conditions for the background wind. The clouds are initialized with a fixed contrast of $100$ in density (or temperature) with the wind. The cloud mass evolution curves in Fig. 7 are colored according to the distance travelled by the cloud (center of mass) with respect to its initial position in the wind, as indicated by the colorbar. As illustrated earlier in Fig. 5, although a simple identification of clouds as cold gas ($T<8\times 10^{4}\ \rm K$) in a CC85 wind shows cloud growth (albeit at a slower rate compared to ‘vanilla’), the clouds fail to remain contrasting features in the wind. This trend holds for clouds of all sizes and occurs when the clouds reach a specific critical distance downstream. Since the cloud density at any instant only depends on its initial density and its distance from the initial position in the wind (see Eq. 3.1 in Section 3.3), the saturation of the ‘contrasting’ cold mass and subsequent turnover from growth to destruction occur at the same distance from the injection region of the wind for all clouds presented here (as they start with the same initial contrast with the background wind). Therefore, the line color in Fig. 7 (denoting the distance travelled by the cloud) at the time of the peak cold+dense mass is the same across simulation parameters (different cloud sizes). The larger clouds with smaller $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}$ and larger $t_{\rm cc,ini}$, travel farther downstream for every $t_{\rm cc,ini}$ and reach the turnover distance in fewer multiples of $t_{\rm cc,ini}$, as seen in Fig. 7 (also see the discussion near the end of Section 3.2).
@@ 3.5 Pressure-density distribution of cold gas
Figure 8: Temporal evolution of the phase space density-pressure distribution of the cold gas ($T<2T_{\rm cl}=8\times 10^{4}\ {\rm K}$) in our fiducial CC85 simulation. The gas density and pressure are normalized by the wind properties at the initial location of the cloud. The different panels are at different times in the simulation as indicated by the blue text annotated in the plots. The gray dashed and dotted lines in all the panels show the isotherms of $T_{\rm cl}$ and $2T_{\rm cl}$, respectively. The orange diamond in the left panel marks the position of the initial cloud material in this phase space, while the red cross in all the panels marks the median density and pressure of the cold gas at the instantaneous position of the center of mass of the cloud (cf. Fig. 6). With time, the cold material becomes low in density. The red cross (center of mass) always lies at the high probability region of the distribution and the distribution itself has a small spread about the $T=T_{\rm cl}$ isotherm due to the elongated nature of the cloud as it moves downstream. Pressure fluctuations $\Delta\mathcal{P}=(p_{\rm cl}-p_{\rm wind})/p_{\rm wind}$ in the cold gas compared to the local wind pressure are shown in the inset plot in each panel. In the inset figures, we plot the mass-weighted distribution of pressure fluctuation in the cold gas $1/M_{0}\ (dM/d\log_{10}\Delta\mathcal{P})$, revealing that the cloud is slightly under-pressurized. An animated version of this plot is available here.
We analyze the distribution of the thermodynamic state of all the cold gas, with temperature less than $2T_{\rm cl}=8\times 10^{4}\ \rm K$. This excludes the wind material that is initially cooler than this threshold, resulting in a smaller sub-domain chosen for the analysis (cf. Section 2.3). We show the distribution function at three different times from one of our cloud-crushing simulations in a CC85 wind with $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}=0.2$ (the fiducial parameters; cf. Table 1) in Fig. 8. As expected, we find that the pressure and density of all gas cells are bound between the two isotherms at $T_{\rm cl}$ and $2T_{\rm cl}$ indicated by dashed lines in the figure. Initially, the cloud is isobaric with the wind, but is $100$ times denser. Therefore, all the cold gas initally occupies only one point in the phase space indicated by the orange diamond in Fig. 8. With time, mixing and radiative cooling cause a spread in the properties of the cold gas, but the median density and temperature at the center of mass of the cold gas, marked by the red cross in the figure, remain very close to the $T_{\rm cl}$ isotherm. As expected, the red cross is a high probability region of the distribution function and can, therefore, be used as a reliable estimate of a representative thermodynamic state for the entire cloud. The inset plots in Fig. 8 show the distribution of the mass-weighted deviation of the pressure in the cold gas compared to the local wind pressure. These plots reveal that the bulk of the cold gas in the cloud attains approximate pressure equilibrium with the local background. Note that there exists a small pressure gradient between the cloud and the wind because the cloud is marginally under-pressurized. This could be either a signature of a local subsonic cooling flow that can potentially siphon cold gas into these clouds (Dutta et al. 2022), especially since such steady solutions exist in cylindrical geometry, or partly a consequence of insufficient resolution to resolve the local cooling length (Fielding et al. 2020; also see Appendix D for a discussion on convergence in our simulations).

### v3

@@ 3.3 Properties of clouds in a typical CC85 wind
In each snapshot from our simulation, starting from the minimum until the maximum radial extent of cold gas along the wind, we choose concentric radial shells in the wind of width $0.25\ R_{\rm cl}$. We calculate the median pressure (mass-weighted), density, and velocity of the cold gas ($T<8\times 10^{4}\ \rm K$) in each of these shells to obtain a cloud profile as a function of the downstream distance (for fiducial parameters; cf. Table 1). These pressure, density, and velocity profiles of the cloud are presented in Fig. 7, where we also show all the cloud cells as an underlying scatter plot to indicate the spread in the data. We use black dashed lines to show the corresponding CC85 wind profile for reference. Different colours are chosen to denote cloud profiles at different times in their evolution, as labelled in the accompanying legends. From these cloud profiles, we find that the cloud pressure, in the left panel of Fig. 7, closely follows the local wind pressure (except for the small portion at the head of the cloud). Therefore, the cloud evolves roughly isobarically in the wind. Note that the median pressure is slightly lower than the local wind pressure which might be responsible for the inflow of cold gas into the cloud as our fiducial parameter corresponds to a growing cloud with strong radiative cooling (see Section 3.5 & Fig. 9 for details). In addition, since the cold gas itself is approximately isothermal (see Fig. 4 & Section 3.5), we can approximate the cloud density as
@@
which implies that $n_{\rm cl}\propto p_{\rm wind}$. Therefore, in the asymptotic limit, $n_{\rm cl}\propto(d_{\rm cl}/d_{\rm cl,ini})^{-2\gamma}$, where $d_{\rm cl}$ and $d_{\rm ini}$ are the distances of the cold gas and the initial position of the cloud respectively from the wind centre. Indeed, the cloud density profile follows our expectation as indicated by the dotted gray curves in the middle panel of Fig. 7. The most important implication is that the cloud loses its density contrast with distance. This is generic for growing clouds of all sizes, causing them to gradually dissolve in the background wind (cf. Section 3.4). The cloud velocity profile in the rightmost panel in Fig. 7 shows that the cloud gradually catches up with the velocity of the wind. As the wind velocity in the CC85 solution is not strictly a constant (but asymptotes to a constant), there is a substantial velocity difference between the head and the tail of the cloud, causing it to stretch in the direction of the wind.
In addition to this, we find that the inflow velocity of gas due to radiative cooling feeding the cold clouds as it moves in the CC85 wind shows only a crude agreement with the theory of turbulent boundary layers introduced by Fielding et al. 2020; Gronke & Oh 2020; Tan et al. 2021 (see Appendix B). Further work is needed to test the limits of applicability of this theory and to come up with better models. The profiles shown in Fig. 7 reveals a consistent picture even at a higher resolution (see Fig. 24) demonstrating the robustness of our scaling relations and our results on the general evolution of cold clouds in an expanding outflow.
@@ 3.4 Clouds as dense cold clumps embedded in a diffuse background
Figure 8: Temporal evolution of mass in clouds, defined as cold gas ($T<2T_{\rm cl}=8\times 10^{4}\ \rm K$) much denser than their local background (more than $10$ times denser than the wind at their location). The gas cells satisfying this criterion stand out as contrasting radial filaments in the wind (cf. Fig. 4). The time is normalised by the initial cloud-crushing time $t_{\rm cc,ini}$ (which is larger for larger and vice versa). Each line in the plot refers to a simulation with a fixed value of $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}$ (inversely proportional to the initial cloud size) – annotated by the numbers adjacent to the curves. The lines are coloured according to the distance the cloud has travelled with respect to its initial position in the wind. Since the Mach number at the initial position $\mathcal{M}=1.496$ of the cloud and its density contrast $\chi=100$ are kept the same across all the simulations presented here, the initial cloud position $d_{\rm cl,ini}$ is the same. The cloud density falls at a steeper rate with distance compared to the the CC85 wind density (cf. Section 3.2). Consequently, the clouds fail to remain cold as dense contrasting features in the wind as they travel downstream. Therefore, after the cloud has travelled a critical distance down the wind (approximately the same in all simulations; cf. Eq. 3.1), the (contrasting) cloud mass transitions from an initially growing phase to a destruction phase. This transition occurs at around $d_{\rm cl}/d_{\rm cl,ini}\sim 4$, where the line colour is blue-cyan (at a time approximately predicted by Eq. 3.3; black crosshair) and the gray dotted and black dashed lines intersect in the middle panel of Fig. 7. Follow this link to see an alternate plot where the quantities for the x-axis and the colourbar are swapped.
Figure 9: Temporal evolution of the phase space density-pressure distribution of the cold gas ($T<2T_{\rm cl}=8\times 10^{4}\ {\rm K}$) in our fiducial CC85 simulation. The gas density and pressure are normalised by the wind properties at the initial location of the cloud. The different panels are at different times in the simulation as indicated by the blue text annotated in the plots. The gray dashed and dotted lines in all the panels show the isotherms of $T_{\rm cl}$ and $2T_{\rm cl}$, respectively. The orange diamond in the left panel marks the position of the initial cloud material in this phase space, while the red cross in all the panels marks the median density and pressure of the cold gas at the instantaneous position of the centre of mass of the cloud (cf. Fig. 7). With time, the cold material becomes low in density. The red cross (centre of mass) always lies at the high probability region of the distribution and the distribution itself has a small spread about the $T=T_{\rm cl}$ isotherm due to the elongated nature of the cloud as it moves downstream. Pressure fluctuations $\Delta\mathcal{P}=(p_{\rm cl}-p_{\rm wind})/p_{\rm wind}$ in the cold gas compared to the local wind pressure are shown in the inset plot in each panel. In the inset figures, we plot the mass-weighted distribution of pressure fluctuation in the cold gas $1/M_{0}\ (dM/d\log_{10}\Delta\mathcal{P})$, revealing that the cloud is slightly under-pressurized. An animated version of this plot is available here.
@@
In the context of cloud-crushing in a CC85 wind, the cloud encounters a cooler adiabatically expanding wind as it moves downstream. Consequently, the cold cloud formed by mixing and subsequent radiative cooling becomes diffuse as it moves further downstream. Therefore, simply identifying all the cold gas as cloud material gives an incomplete picture of the ongoing process. Considering this, we adopt a more restrictive definition of cloud, identifying it as cold and dense features (rather than just cold) formed via mixing and radiative cooling, relative to the featureless expanding wind (cf. Section 2.3). In the remainder of this section we illustrate that in the presence of a CC85 wind, the clouds in the cooling regime grow in their cold mass but at a reduced rate (see Section 3.2 & Fig. 5), and in addition continue to expand compared to clouds in a uniform plane-parallel wind. It is worth noting that the ratio of the contrast (with the background) that a cloud has between any two distances down the wind is the same for clouds of all sizes (cf. Section 3.2) if the clouds start from the same initial position with the same contrast. This ensures that the qualitative picture that clouds lose contrast while moving downstream remains unaltered by the choice of threshold density contrast ($10$ here; see Section 2.3).
@@
We illustrate the evolution of the cloud mass identified using our ‘contrasting cloud identification’ criterion (discussed in Section 2.3) in Fig. 8. The different curves in the figure correspond to different initial cloud sizes in our simulation, labelled by $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}$. The clouds in different runs from our simulation presented here are initially placed at the same position with respect to the wind (for same Mach number $\mathcal{M}_{\rm ini}$) with identical conditions for the background wind. The clouds are initialised with a fixed contrast of $100$ in density (or temperature) with the wind. The cloud mass evolution curves in Fig. 8 are coloured according to the distance travelled by the cloud (centre of mass) with respect to its initial position in the wind, as indicated by the colourbar. As illustrated earlier in Fig. 5, although a simple identification of clouds as cold gas ($T<8\times 10^{4}\ \rm K$) in a CC85 wind shows cloud growth (albeit at a slower rate compared to ‘vanilla’), the clouds fail to remain contrasting features in the wind. This trend holds for clouds of all sizes and occurs when the clouds reach a specific critical distance downstream. Since the cloud density at any instant only depends on its initial density and its distance from the initial position in the wind (see Eq. 3.1 in Section 3.2), the saturation of the ‘contrasting’ cold mass and subsequent turnover from growth to destruction/fading occur at the same distance from the injection region of the wind for all clouds presented here (as they start with the same initial contrast with the background wind). Therefore, the line colour in Fig. 8 (denoting the distance travelled by the cloud) at the time of the peak cold+dense mass is the same across simulation parameters (different cloud sizes). The larger clouds, with smaller $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}$ and larger $t_{\rm cc,ini}$, travel farther downstream for every $t_{\rm cc,ini}$ and reach the turnover distance in fewer multiples of $t_{\rm cc,ini}$, as seen in Fig. 8 (also see the discussion near the end of Section 3.2).
The reduction of ‘contrasting’ cold mass constitutes a physically relevant as well as observationally detectable state in the evolution of cold clouds already growing in an expanding wind. Therefore, we term this as a ‘fade-away’ regime in the evolution of growing clouds. It is worth noting that this is different from the usual growth/destruction regime of cold gas, which depends only on the radiative cooling of the mixed boundary layer around the cloud. As illustrated, clouds, which are initially growing, gradually transitions to a ‘fade-away’ regime as they move down an expanding wind. We make an empirical estimate of the time when the mass in cold+dense clouds transition from their growing state to the ‘fade-away’ state in a CC85 wind. As the cloud moves downstream, its temperature remains close to $T_{\rm cl}$ while the wind temperature $T_{\rm wind}$ scales as $d^{-2(\gamma-1)}$, where $d$ is the radial distance of the cloud from the wind centre. Since we adopt a cut-off density contrast of $10$, the transition from growth to fading is expected at a fixed position $d_{f}$ down the wind when the density contrast drops from an initial contrast of $\chi_{\rm ini}=100$ (in our simulation) to $\chi_{f}=10$ (our adopted cut-off). We propose an empirical estimate of the time elapsed $t_{\rm turn}$ when the cloud has travelled this distance downstream. The distance a cloud travels downstream starting from its initial position $d_{f}-d_{\rm ini}$ can be approximated as
$d_{f}-d_{\rm ini}=\frac{1}{2}\frac{v_{\rm wind,\infty}}{f_{\rm acc}t_{\rm drag,ini}}t_{\rm turn}^{2},$ (3.2)
where we assume that the cloud accelerates uniformly to the asymptotic wind speed $v_{\rm wind,\infty}$ ($\sqrt{2\dot{E}/\dot{M}}$; $\dot{E}$ and $\dot{M}$ are energy and mass injection rates in the wind) in some multiple $f_{\rm acc}\gtrsim 1$ of the initial drag time $t_{\rm drag,ini}$. Further, using the scaling of the density contrast $\chi_{f}/\chi_{\rm ini}=(d_{f}/d_{\rm ini})^{-2(\gamma-1)}$ we obtain,
$\frac{t_{\rm turn}}{t_{\rm cc,ini}}=\sqrt{2f_{\rm acc}\left(\frac{d_{\rm ini}}{R_{\rm cl}}\right)\left[\left(\frac{\chi_{\rm ini}}{\chi_{f}}\right)^{\frac{1}{2(\gamma-1)}}-1\right]}.$ (3.3)
Appropriate to our definition of the threshold over-density contrast $\chi_{f}=10$, we find that $f_{\rm acc}=1.4$ approximately (within a factor of a few) explains our simulation results in estimating the time when ‘contrasting’ cloud mass stops growing in Fig. 8.
@@ 3.5 Pressure-density distribution of cold gas
We analyse the distribution of the thermodynamic state of all the cold gas, with temperature less than $2T_{\rm cl}=8\times 10^{4}\ \rm K$. This excludes the wind material that is initially cooler than this threshold, resulting in a smaller subdomain chosen for the analysis (cf. Section 2.3). We show the distribution function at three different times from one of our cloud-crushing simulations in a CC85 wind with $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}=0.2$ (the fiducial parameters; cf. Table 1) in Fig. 9. As expected, we find that the pressure and density of all gas cells are bound between the two isotherms at $T_{\rm cl}$ and $2T_{\rm cl}$ indicated by dashed lines in the figure. Initially, the cloud is isobaric with the wind, but is $100$ times denser. Therefore, all the cold gas initially occupies only one point in the phase space indicated by the orange diamond in Fig. 9. With time, mixing and radiative cooling cause a spread in the properties of the cold gas, but the median density and temperature at the centre of mass of the cold gas, marked by the red cross in the figure, remain very close to the $T_{\rm cl}$ isotherm. As expected, the red cross is a high probability region of the distribution function and can, therefore, be used as a reliable estimate of a representative thermodynamic state for the entire cloud. The inset plots in Fig. 9 show the distribution of the mass-weighted deviation of the pressure in the cold gas compared to the local wind pressure. These plots reveal that the bulk of the cold gas in the cloud attains approximate pressure equilibrium with the local background. Note that there exists a small pressure gradient between the cloud and the wind because the cloud is marginally under-pressurized. This could be either a signature of a local subsonic cooling flow that can potentially siphon cold gas into these clouds (Dutta et al. 2022), especially since such steady solutions exist in cylindrical geometry and turbulent transport (Tan & Fielding 2024) or partly a consequence of insufficient resolution to resolve the local cooling length (Fielding et al. 2020). Such under-pressurized clouds persist even at higher resolution (cf. Fig. 24), making it less likely to be a simple resolution artifact.